Consider the proton–proton cycle that occurs in most stars (including our own Sun): Step 1: 1H + 1H → 2H + e+ + νe Step 2: 2H + 1H → 3He + γ Step 3: 3He + 3He → 4He + 2 1H + γ Calculate the net energy released from the three steps. Do not ignore the mass of the positron in Step 1. (You may ignore the mass of the neutrino.) MeV
Answer #1

Step 1: 1H + 1H → 2H + e+ + \nu + energy

Step 2: 2H + 1H → 3He + γ + energy

Step 3: 3He + 3He → 4He + 21H + γ + energy

So Net reaction is ( step1 x 2 + step2 x 2 + step3)

41H →4He + 2 e++ 2 ν + 3 γ + energy

Mass of a proton is 1.6726 × 10−27kg , So the mass of 4 protons = 6.6904 × 10−27kg

The mass of a 4He nucleus is 6.6447 × 10−27kg

The mass of positron is 9.1094×10−31 kg , Somass of 2 positron = 18.2188 x 10-31 kg = 0.0018 x 10-27 kg

Missing mass(\Delta m) = Mass of 4 protons - (Mass of 4He + Mass of 2 positron) =

                                6.6904 × 10−27 - (6.6447 × 10−27 + 0.0018 x 10-27 ) = 0.0439x10-27 kg

The missing mass was converted into energy by Einstein’s famous equation
E = \Deltamc2 = 0.0439x10-27 x (3 x 108 )2 = 3.951 x 10-12 J

1J = 6.242 x 1012 Mev

Hence 3.951 x 10-12 = 24.6602Mev

please rate!